The surface area of an average naked human adult is 2.2 m^2 and the average skin temperature is 33 degrees C. The emissivity of human skin is 0.95. Determine the radiation heat transfer rate from the human to the walls of a room if they are at 22 degrees C. This is why naked person feels chilly at room temperature.
Answer #1

T= 33C @ 306K Griver A=227?, &= 0-95, T2 = 22°C (or) 295k We have O= Boltzzman Constant = 5.68 x 10 W mukh Q = 0-95-56340%-22